OpenStudy (anonymous):

\[2\sqrt{18} \times 7\sqrt{6}\] can someone check my work radicals?

OpenStudy (anonymous):

\[2\sqrt{18} \times 7\sqrt{6}\]

OpenStudy (anonymous):

show your work and I will check it

OpenStudy (anonymous):

is the answer by any chance \[84\sqrt{3}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh thank you!

OpenStudy (anonymous):

\[-3\sqrt{20} \times \sqrt{5}\]

OpenStudy (anonymous):

15 not 5

OpenStudy (anonymous):

\[-30\sqrt{3}\] maybe?